OpenStudy (anonymous):

f(x) = 4x + 6 g(x) = 2x + 10 At what input would both functions produce the same output?

OpenStudy (anonymous):

equate them and solve for x: 4x+6=2x+10

OpenStudy (anonymous):

10-6?

OpenStudy (anonymous):

2x=10-6 2x=4 x=2

OpenStudy (anonymous):

oh ok thank youu